# B3C: A Minimalist Approach to Offline Multi-Agent Reinforcement Learning

### v2

@@ 4.2 Numerical Results
Figure 5: Ablation study on the clipping value, $M$: Performance of FACMAC+B3C with respect to $M$. $M=\infty$ (no clip) corresponds to FACMAC+BC. The performance of OMIGA is also included as a blue dotted line. Error bars represent one standard deviation.
Figure 6: Ablation study on value factorization: Performance differences of VDN and monotonic factorization compared to non-monotonic factorization.
4.3 Analysis
To gain deeper insights into the proposed method, B3C, and value factorization techniques for offline learning, we present several ablation studies in the following sections.
4.3.1 Critic Clipping: B3C versus BC
Critic clipping alleviates overestimation by directly constraining the target value in policy evaluation. As mentioned earlier, BC regularization alone often makes learning unstable, and even the critic diverges. We have already shown instability problem in Sec. 4.2 through the worst performance of MA-TD3+BC in the multi-agent particle environments. Here, we provide further analysis of how critic clipping addresses this issue for better understanding in multi-agent Mujoco environments. Fig. 3 presents the test return and the target value, $y^{jt}$ from Eq. 3.2, during the training of FACMAC+BC and FACMAC+B3C. The results indicate that FACMAC+BC, which relies solely on BC regularization, results in diverging value estimates. When this divergence occurs, marked by the dotted line in Fig. 3, the performance correspondingly degrades. However, B3C effectively mitigates this divergence—it occasionally tends to overestimate but either stabilizes quickly or avoids divergence entirely, ensuring more consistent performance. Note that this phenomenon does not always occur but is more likely when the RL objective is weighted more heavily than BC regularization; however, critic clipping can prevent it and enable stable learning.
@@ 4.3.1 Critic Clipping: B3C versus BC
4.3.2 Critic Clipping: Clipping Value
@@
CC clips the Q-function in the target value by the scaled maximum return in the dataset. This introduces a hyperparameter—the clipping value, $M$ in Eq. 3.2. As discussed in Sec. 3.2, $M=1$ is sufficient since the maximum return in the dataset is already high enough to handle overestimation. We conduct an ablation study regarding the clipping value. Fig. 5 shows the performance of FACMAC+B3C by varying $M$ across $1/10$, $1/4$, $1$, and $4$. Additionally, we include the performance of FACMAC+BC, which does not use critic clipping, as $M=\infty$, and OMIGA (Wang et al., 2024) as a blue dotted line. It is seen that $M=1$ achieves the best performance in all cases except for the HC-k1-e-m2 environment. While extremely small or large values of $M$ perform worse than $M=1$, FACMAC+B3C with any of the tested $M$ values still outperforms both FACMAC+BC and the SOTA algorithm, OMIGA. In the HC-k1-e-m2 environment, where FACMAC+B3C performs suboptimally compared to the dataset quality, reducing $M$ improves performance. Overall, $M=1$ is a reasonable choice, providing the best performance while minimizing the burden of hyperparameter tuning.
4.3.3 Value factorization
@@
We here provide an ablation study of the empirical observation described in Sec. 3.3 that non-monotonic factorization is better than monotonic factorization in offline settings. Fig. 6 shows the performance differences of VDN and monotonic factorization compared to non-monotonic factorization. It is seen that non-mono generally performs better than both mono and vdn, though linear factorization shows better performance in certain cases. However, mono performs poorly in offline settings, which differs from the online settings (Peng et al., 2021).
5 Conclusion
@@
In this paper, we have proposed a simple regularization technique named behavior cloning with critic clipping (B3C) for offline multi-agent RL that addresses overestimation and over-regularization while ensuring stability. In addition, we have investigated existing value factorization techniques in offline settings, which are understudied, providing the observation that non-monotonic factorization performs better than monotonic factorization. Numerical results show that B3C with non-monotonic factorization yields outstanding performance across the considered environments with diverse dataset levels. We have also provided several analyses and ablation studies to illustrate how the proposed method operates and affects learning.
Impact Statement
Potential societal impacts of this research may arise in areas where multi-agent systems play a critical role, such as autonomous vehicles, distributed resource management, and multi-robot collaboration. Our approach contributes to safer and more reliable decision-making processes in these systems, thereby promoting ethical AI development.
Appendix A Hyperparameters
@@
We here provide the hyperparameters used for the proposed method. We have three hyperparameters: the RL coefficient $\alpha$, the BC coefficient $\beta$, and the clipping value $M$. As discussed in the main paper, we set $M=1$ except for one task (Ant, medium-replay dataset). For multi-agent particle environments, the values for $\alpha$ and $\beta$ are provided in Table 3. We use $M=1$ for all cases. Next, for multi-agent Mujoco environments, we fix $\beta=1$. The values for $\alpha$ and $M$ are provided in Table. 4.
Table 3: Hyperparameters in multi-agent particle environments
@@
Table 4: Hyperparameters for Various Environments
Environment Dataset $\alpha$ $M$
Hopper expert 16 1
medium 8 1
medium-replay 0.25 0.1
medium-expert 1 1
Ant expert 1 1
medium 1 1
medium-replay 1 1
medium-expert 0.5 1
HalfCheetah expert 16 1
medium 16 1
medium-replay 16 1
medium-expert 16 1
HC_obsk0 expert 1 1
medium1 1 1
medium2 4 1
expert-medium1 1 1
expert-medium2 4 1
medium1-medium2 4 1
HC_obsk1 expert 1 1
medium1 1 1
medium2 4 1
expert-medium1 4 1
expert-medium2 1 1
medium1-medium2 4 1
Swimmer expert 1 1
medium1 1 1
medium2 1 1
expert-medium1 4 1
expert-medium2 1 1
medium1-medium2 0.5 1
@@
Figure 7: Performance difference FACMAC+BC compared to FACMAC+B3C, with negative values indicating that FACMAC+BC performs worse than FACMAC+B3C. A value of $-5\%$ represents that BC performs $5\%$ worse than B3C.

### v3

@@ 4.3.1. Critic Clipping: B3C versus BC
Figure 6. Ablation study on the clipping value, $M$: Performance of FACMAC+B3C with respect to $M$. $M=\infty$ (no clip) corresponds to FACMAC+BC. The performance of OMIGA is also included as a blue dotted line. Error bars represent one standard deviation.
Figure 7. Ablation study on value factorization: Performance differences of VDN and monotonic factorization compared to non-monotonic factorization.
@@
4.3.2. Critic Clipping: Clipping Value
@@
CC clips the Q-function in the target value by the scaled maximum return in the dataset. This introduces a hyperparameter—the clipping value, $M$ in Eq. 3.2. As discussed in Sec. 3.2, $M=1$ is sufficient since the maximum return in the dataset is already high enough to handle overestimation. We conduct an ablation study regarding the clipping value. Fig. 6 shows the performance of FACMAC+B3C by varying $M$ across $1/10$, $1/4$, 11, and $4$. Additionally, we include the performance of FACMAC+BC, which does not use critic clipping, as $M=\infty$, and OMIGA wang2024offline as a blue dotted line. It is seen that $M=1$ achieves the best performance in all cases except for the HC-k1-e-m2 environment. While extremely small or large values of $M$ perform worse than $M=1$, FACMAC+B3C with any of the tested $M$ values still outperforms both FACMAC+BC and the SOTA algorithm, OMIGA. In the HC-k1-e-m2 environment, where FACMAC+B3C performs suboptimally compared to the dataset quality, reducing $M$ improves performance. Overall, $M=1$ is a reasonable choice, providing the best performance while minimizing the burden of hyperparameter tuning.
4.3.3. Value factorization
@@
We here provide an ablation study of the empirical observation described in Sec. 3.3 that non-monotonic factorization performs better than monotonic factorization in offline settings. Fig. 7 shows the performance differences of VDN and monotonic factorization compared to non-monotonic factorization across all environments. Overall, non-mono consistently outperforms both mono and vdn, though the linear factorization occasionally achieves comparable or slightly higher scores in certain easy datasets.
This result contrasts with prior findings in online settings peng2021facmac, where monotonic factorization often yielded stable and competitive performance. We attribute this difference to the distributional shift inherent in offline learning: enforcing monotonicity restricts the representational capacity of the critic, making it difficult to approximate complex joint action dependencies when unseen samples dominate the dataset. In contrast, non-monotonic factorization allows the critic to flexibly capture nonlinear inter-agent correlations, leading to more accurate value estimation and better policy learning.
5. Conclusion
@@
In this paper, we have proposed a simple regularization technique named behavior cloning with critic clipping for offline multi-agent RL that addresses overestimation and over-regularization while ensuring stability. In addition, we have investigated existing value factorization techniques in offline settings, providing the observation that non-monotonic factorization performs better than monotonic factorization. Numerical results show that B3C with non-monotonic factorization yields outstanding performance across the considered environments with diverse dataset levels. We have also provided several analyses and ablation studies to illustrate how the proposed method operates and affects learning.
Limitation Our work lacks theoretical analysis and is empirically driven. We consider this adequate for our minimalist objective and leave theoretical extensions for future work.
{acks}
This work was supported by the DARPA EMHAT Program under Agreement No. HR00112490409 and by the ONR Award No. N00014-23-1-2840.
Appendix A Training Details
We implemented FACMAC+B3C based on the public code of OMIGA wang2024offline, and conducted the experiments on a server with dual AMD EPYC 7713 CPUs and an NVIDIA RTX 6000 Ada Generation GPU. Each experiement took about 4 hours. Code and experiment scripts are available at: https://github.com/wjkim1202/B3C_OfflineMARL
All policy and critic networks are implemented as multi-layer perceptrons (MLPs) with hidden layer sizes of 256. We use the Adam optimizer with a learning rate of $5\times 10^{-4}$, a discount factor $\gamma=0.99$, and Polyak target updates with $\tau=0.005$. Gradients are clipped to a maximum norm of 1.0, and training is performed with a batch size of 128.
@@ Appendix A Training Details
We now provide the key hyperparameters used for the proposed method. We have three key hyperparameters: the RL coefficient $\alpha$, the BC coefficient $\beta$, and the clipping value $M$. As discussed in the main paper, we set $M=1$ except for one task (Ant, medium-replay dataset). For multi-agent particle environments, the values for $\alpha$ and $\beta$ are provided in Table 4. We use $M=1$ for all cases. Next, for multi-agent Mujoco environments, we fix $\beta=1$. The values for $\alpha$ and $M$ are provided in Table. 3.
Table 3. Hyperparameters for Various Environments
Environment Dataset $\alpha$ $M$
Hopper expert 16 1
medium 8 1
medium-replay 0.25 0.1
medium-expert 1 1
Ant expert 1 1
medium 1 1
medium-replay 1 1
medium-expert 0.5 1
HalfCheetah expert 16 1
medium 16 1
medium-replay 16 1
medium-expert 16 1
HC_obsk0 expert 1 1
medium1 1 1
medium2 4 1
expert-medium1 1 1
expert-medium2 4 1
medium1-medium2 4 1
HC_obsk1 expert 1 1
medium1 1 1
medium2 4 1
expert-medium1 4 1
expert-medium2 1 1
medium1-medium2 4 1
Swimmer expert 1 1
medium1 1 1
medium2 1 1
expert-medium1 4 1
expert-medium2 1 1
medium1-medium2 0.5 1
Table 4. Hyperparameters in multi-agent particle environments
@@
Figure 8. Performance difference FACMAC+BC compared to FACMAC+B3C, with negative values indicating that FACMAC+BC performs worse than FACMAC+B3C. A value of $5\%$ represents that BC performs $5\%$ worse than B3C.